# Boundedness of commutator generated by fractional integral operator and Orlicz-BMO function

### 25 Apr 2026 v1

@@ 1. Introduction
Two classical results (see [9, Theorem 1.2.3] and [11, Theorem 4]) asserted that there exists constants $C$ such that
@@
with $1<p<q<\infty$ and $1/q-1/p=\alpha/n$. In the framework of Hardy spaces, the boundedness of $I_{\alpha}$ is considered by Stein and Weiss [21, Theorem G] in their pioneering work on multi-variable harmonic functions, they established that
@@
where $g$ is a continuous function whose support is compact. A famous theorem of Chanillo [3] characterizes the $\mathrm{BMO}({{{\mathbb{R}}}^{n}})$ space via the boundedness of $b,I_{\alpha}]$, that is, $b\in\mathrm{BMO}({{{\mathbb{R}}}^{n}})$ if and only if the commutator $b,I_{\alpha}]$ is bounded from $L^{p}({{{\mathbb{R}}}^{n}})$ to $L^{q}({{{\mathbb{R}}}^{n}})$, where $0<\alpha<n$ and $1<p<\frac{n}{\alpha}$ with $1/q=1/p-\alpha/n$. Paluszyński [19, Theorem 1.1] proved that $b,I_{\alpha}]$ does not map from $H^{1}({{{\mathbb{R}}}^{n}})$ into $L^{\frac{n}{n-\alpha}}({{{\mathbb{R}}}^{n}})$ continuously. Later, Pérez [20] introduced a subspace $\mathcal{H}^{1}_{b}({{{\mathbb{R}}}^{n}})\subset H^{1}({{{\mathbb{R}}}^{n}})$ in the study of commutator of Calderón-Zygmund operator. Prompted by the bilinear splitting technique applied to the pair of Hardy space $H^{1}({{{\mathbb{R}}}^{n}})$ and $\mathrm{BMO}$ space, which is originally introduced by Bonami and Grellier [1], an important work is that Ky [13] found a suitable subspace $H_{b}^{1}({{{\mathbb{R}}}^{n}})$ of $H^{1}({{{\mathbb{R}}}^{n}})$ and established that $b,I_{\alpha}]$ have boundedness from $H_{b}^{1}({{{\mathbb{R}}}^{n}})$ to $L^{\frac{n}{n-\alpha}}({{{\mathbb{R}}}^{n}})$. Moreover, Ky [13] also proved that the $b,I_{\alpha}]$ maps $H^{1}({{{\mathbb{R}}}^{n}})$ continuously into the weak Lebesgue space $L^{\frac{n}{n-\alpha},\,\infty}({{{\mathbb{R}}}^{n}})$. For more related results, we refer to the interested reader to [3, 4, 10, 15, 20].
@@
If $\varphi(t)=t$, $\mathrm{BMO}_{\varphi}(\mathbb{R}^{n})=\mathrm{BMO}(\mathbb{R}^{n})$. For more research on $\mathrm{BMO}$ type spaces, see [6, 7, 8, 22].
@@
Theorems 1.5 and 1.6 are a further extension for the results of Ky [13, Corollary 8.1 and Theorem 8.2] in the Orlicz-type spaces framework. A study related to the Calderón-Zygmund operator, see [5].
@@ 2. Class $\mathcal{K}_{\varphi,\,\alpha}$ and typical example
To facilitate subsequent proof, it is proposed that the result in [5, (2.3)] be employed as a lemma.
@@
Accoording to[17, Theorem 3.1] and [9, Theorem 1.2.3], $I_{\alpha}$ has boundedness from $H^{1}(\mathbb{R}^{n})$ to $L^{\frac{n}{n-\alpha}}({{{\mathbb{R}}}^{n}})$, and $I_{\alpha}$ also bound from $L^{1}(\mathbb{R}^{n})$ to $L^{\frac{n}{n-\alpha},\,\infty}({{{\mathbb{R}}}^{n}})$. In order to prove this lemma, it suffices to show that $I_{\alpha}$ satisfies Definition 2.3(iii).
@@
To estimate $\rm I$}, if $q\in(1,n/\alpha)$ and $1/q_{0}=1/q-\alpha/n$, then we see that $I_{\alpha}$ is bounded from $L^{q}(\mathbb{R}^{n})$ to $L^{q_{0}}(\mathbb{R}^{n})$. Taking $r>1$ with $1/r+1/q_{0}=1/p_{\alpha}$, by Hölder inequality, Lemma 2.4 and [5, Lemma 3.4 and Lemma 5.5], it follows that
@@
This, along with lemma 2.4 and [5, Lemma 3.4 and Lemma 5.5], yields
@@ 3. Decomposition of commutator
Note that (i) and (ii) follow from [5, Propositions 3.14, 3.16 and 3.17]. By [5, (5.17)], we obtain that (iii) holds. ∎
@@
Given a growth function $\varphi$ and its lower and upper critical indices satisfy (1.4) and (1.5). Let $\Pi_{4}$ be defined as (3.15) and $T\in K_{\varphi,\,\alpha}$ with $0<\alpha<n$. For any $b\in\mathrm{BMO}_{\varphi}(\mathbb{R}^{n})$ and any $g\in H^{\varphi}(\mathbb{R}^{n})$, there exists an operator
@@
Based on the argument of [16, Proposition 6.7], the following three steps will be used to prove the desired result.
Step 1: Verifying (3.16) under the case that $g$ has a finite linear combination of wavelets. We unitize [5, Lemma 3.3, (c) and (a)] to obtain that $g$ has finite atomic decomposition
@@
where $\operatorname{supp}a_{k}\subseteq R_{k}$ and $\langle g,\psi^{\lambda}_{I}\rangle$ is the coefficients of $a_{k}$. Further, let $B_{k}$ be the smallest ball containing dyadic cube $R_{k}$. According to [5, Proposition 2.5], we have
@@
For almost every $x\in\mathbb{R}^{n}$, using a standard algebraic identity (see [16, (6.9)]), it holds
@@
By this, $T\in K_{\varphi,\,\alpha}$, Proposition 3.1(iii) , (3.17) and $\|a_{k}\|_{L^{q}_{B_{k}}}\leq\|a_{k}\|_{L^{\infty}_{B_{k}}}$ with $q\in(1,n/\alpha)$, we see that
@@
Step 3: The proof of (3.16). Based on the facts obtained in Steps 1 and 2, we repeat the argument of [5, (5.18)] to obtain (3.16). This implies Lemma 3.2. ∎
@@
The primary objective of this section is to complete the proofs of Theorems 1.5 and 1.6 by using Lemma 4.1. We first recall the following lemma from [5, Theorem 5.10].

### 28 Apr 2026 v2

@@ 1. Introduction
Two classical results (see [10, Theorem 1.2.3] and [12, Theorem 4]) asserted that there exists constants $C$ such that
@@
with $1<p<q<\infty$ and $1/q-1/p=\alpha/n$. In the framework of Hardy spaces, the boundedness of $I_{\alpha}$ is considered by Stein and Weiss [26, Theorem G] in their pioneering work on multi-variable harmonic functions, they established that
@@
where $g$ is a continuous function whose support is compact. A famous theorem of Chanillo [4] characterizes the $\mathrm{BMO}({{{\mathbb{R}}}^{n}})$ space via the boundedness of $b,I_{\alpha}]$, that is, $b\in\mathrm{BMO}({{{\mathbb{R}}}^{n}})$ if and only if the commutator $b,I_{\alpha}]$ is bounded from $L^{p}({{{\mathbb{R}}}^{n}})$ to $L^{q}({{{\mathbb{R}}}^{n}})$, where $0<\alpha<n$ and $1<p<\frac{n}{\alpha}$ with $1/q=1/p-\alpha/n$. Paluszyński [22, Theorem 1.1] proved that $b,I_{\alpha}]$ does not map from $H^{1}({{{\mathbb{R}}}^{n}})$ into $L^{\frac{n}{n-\alpha}}({{{\mathbb{R}}}^{n}})$ continuously. Later, Pérez [23] introduced a subspace $\mathcal{H}^{1}_{b}({{{\mathbb{R}}}^{n}})\subset H^{1}({{{\mathbb{R}}}^{n}})$ in the study of commutator of Calderón-Zygmund operator. Prompted by the bilinear splitting technique applied to the pair of Hardy space $H^{1}({{{\mathbb{R}}}^{n}})$ and $\mathrm{BMO}$ space, which is originally introduced by Bonami and Grellier [2], an important work is that Ky [14] found a suitable subspace $H_{b}^{1}({{{\mathbb{R}}}^{n}})$ of $H^{1}({{{\mathbb{R}}}^{n}})$ and established that $b,I_{\alpha}]$ have boundedness from $H_{b}^{1}({{{\mathbb{R}}}^{n}})$ to $L^{\frac{n}{n-\alpha}}({{{\mathbb{R}}}^{n}})$. Moreover, Ky [14] also proved that the $b,I_{\alpha}]$ maps $H^{1}({{{\mathbb{R}}}^{n}})$ continuously into the weak Lebesgue space $L^{\frac{n}{n-\alpha},\,\infty}({{{\mathbb{R}}}^{n}})$. For more related results, we refer to the interested reader to [1, 5, 11, 13, 16, 21, 24, 25].
@@
If $\varphi(t)=t$, $\mathrm{BMO}_{\varphi}(\mathbb{R}^{n})=\mathrm{BMO}(\mathbb{R}^{n})$. For more research on $\mathrm{BMO}$ type spaces, see [1, 7, 8, 9, 27].
@@
Theorems 1.5 and 1.6 are a further extension for the results of Ky [14, Corollary 8.1 and Theorem 8.2] in the Orlicz-type spaces framework. A study related to the Calderón-Zygmund operator, see [6].
@@ 2. Class $\mathcal{K}_{\varphi,\,\alpha}$ and typical example
To facilitate subsequent proof, it is proposed that the result in [6, (2.3)] be employed as a lemma.
@@
Accoording to[19, Theorem 3.1] and [10, Theorem 1.2.3], $I_{\alpha}$ has boundedness from $H^{1}(\mathbb{R}^{n})$ to $L^{\frac{n}{n-\alpha}}({{{\mathbb{R}}}^{n}})$, and $I_{\alpha}$ also bounded from $L^{1}(\mathbb{R}^{n})$ to $L^{\frac{n}{n-\alpha},\,\infty}({{{\mathbb{R}}}^{n}})$. In order to prove this lemma, it suffices to show that $I_{\alpha}$ satisfies Definition 2.3(iii).
@@
To estimate $\rm I$}, if $q\in(1,n/\alpha)$ and $1/q_{0}=1/q-\alpha/n$, then we see that $I_{\alpha}$ is bounded from $L^{q}(\mathbb{R}^{n})$ to $L^{q_{0}}(\mathbb{R}^{n})$. Taking $r>1$ with $1/r+1/q_{0}=1/p_{\alpha}$, by Hölder inequality, Lemma 2.4 and [6, Lemma 3.4 and Lemma 5.5], it follows that
@@
This, along with lemma 2.4 and [6, Lemma 3.4 and Lemma 5.5], yields
@@ 3. Decomposition of commutator
Note that (i) and (ii) follow from [6, Propositions 3.14, 3.16 and 3.17]. By [6, (5.17)], we obtain that (iii) holds. ∎
@@
Given a growth function $\varphi$ and its lower and upper critical indices satisfy (1.4) and (1.5). Let $\Pi_{4}$ be defined as (3.15) and $T\in\mathcal{K}_{\varphi,\,\alpha}$ with $0<\alpha<n$. For any $b\in\mathrm{BMO}_{\varphi}(\mathbb{R}^{n})$ and any $g\in H^{\varphi}(\mathbb{R}^{n})$, there exists an operator
@@
Based on the argument of [18, Proposition 6.7], the following three steps will be used to prove the desired result.
Step 1: Verifying (3.16) under the case that $g$ has a finite linear combination of wavelets. We unitize [6, Lemma 3.3, (c) and (a)] to obtain that $g$ has finite atomic decomposition
@@
where $\operatorname{supp}a_{k}\subseteq R_{k}$ and $\langle g,\psi^{\lambda}_{I}\rangle$ is the coefficients of $a_{k}$. Further, let $B_{k}$ be the smallest ball containing dyadic cube $R_{k}$. According to [6, Proposition 2.5], we have
@@
For almost every $x\in\mathbb{R}^{n}$, using a standard algebraic identity (see [18, (6.9)]), it holds
@@
By this, $T\in\mathcal{K}_{\varphi,\,\alpha}$, Proposition 3.1(iii) , (3.17) and $\|a_{k}\|_{L^{q}_{B_{k}}}\leq\|a_{k}\|_{L^{\infty}_{B_{k}}}$ with $q\in(1,n/\alpha)$, we see that
@@
Step 3: The proof of (3.16). Based on the facts obtained in Steps 1 and 2, we repeat the argument of [6, (5.18)] to obtain (3.16). This implies Lemma 3.2. ∎
@@
The primary objective of this section is to complete the proofs of Theorems 1.5 and 1.6 by using Lemma 4.1. We first recall the following lemma from [6, Theorem 5.10].